OpenStudy (anonymous):

Solve 16x^4 - 16x^3 - 32x^2 + 36x - 9 = 0.

ganeshie8 (ganeshie8):

factor first 3 terms separately, and last two terms separately

ganeshie8 (ganeshie8):

16x^4 - 16x^3 - 32x^2 + 36x - 9 = 0 --------------------- -------

OpenStudy (anonymous):

how would i factor 36x-9

ganeshie8 (ganeshie8):

we can pull out 9

ganeshie8 (ganeshie8):

36x - 9 9*4x - 9 9(4x-1)

ganeshie8 (ganeshie8):

like that :)

ganeshie8 (ganeshie8):

just try to factor out the common thing from first 3 terms also

OpenStudy (anonymous):

16x^2(x+1)(x-2)

OpenStudy (anonymous):

is the first 3

ganeshie8 (ganeshie8):

yes you're right. but i see we're stuck :| wat we did is of no use. let me think a bit

OpenStudy (anonymous):

oh okay!

ganeshie8 (ganeshie8):

im not getting any ideas sorry :( @hartnn if u have some time...

OpenStudy (anonymous):

it's okay, thanks for trying to help!

hartnn (hartnn):

this is not easy, it may become easy if you know one root/zero

OpenStudy (anonymous):

I don't know any of the roots or zeros.

OpenStudy (anonymous):

my options are a. +-1/2, +-3/2 b. 1/2, +-3/2 c. +-1/2, 3/2 d. +-3/2, 0

hartnn (hartnn):

omg , you have options!! that makes it so simple :)

hartnn (hartnn):

then whats stopping you from just plugging in those numbers one by one and see which one actually satisfies the original equation ?

OpenStudy (anonymous):

I just didn't know if there was a certain equation I should use to solve problems like this or not.

hartnn (hartnn):

put x= 1/2 in 16x^4 - 16x^3 - 32x^2 + 36x - 9 and check whether you get it =0 or not.

OpenStudy (anonymous):

okay. I just didn't know if there was a way other than plugging in all the options. Thanks.

hartnn (hartnn):

there might be a way, but its not easy to think, not even for me. so the best thing to do is to check options :)